I need good explanation on the workings that easier to follow
. A1 Normal No Spac.. Heding1 HeadingT Paragraph Styles 10-1 11 12 113 14-115. Börieis elling ury hamburgers withgold sprayedon top and the demand function for his foodisq-IS2P. What isthe price elasticity ofden and when price is 10?-Nae
Answer #1

Elasticity of demand = Percentage change in quantity demanded / Percentage change in price

= P/Q x dQ/dP

At P = 10,

Q = 15 - 2(10) = 15 - 20 = - 5

dQ/dP = - 2

Ed = 10/(-5) x - 2 = 4